Two smooth spheres P,Q each of radius 25 cm and weighing 500N, rests in a horizontal channel having vertical walls. if the distance between the walls is 90 cm, make calculations for the pressure exerted on the wall and floor at points A,B and C.
1
Expert's answer
2020-09-14T10:19:17-0400

solution

given data

radius of sphere P and Q =25cm

weight of sphere P and Q =500N

diagram for this question can be drawn as





whole system is in equilibrium so all force in x direction and y direction are equal

contact force at point C


"N_c=\frac{w}{tan\theta}=\frac{500\times4}{3}=\frac{2000}{3}N"

contact force at contact of sphere

"N=\frac{w}{sin\theta}=\frac{500\times5}{3}=\frac{2500}3N"


contact force at point A


"N_A=Ncos\theta =\frac{2000}{3}N"


contact force at point B


"N_B=Nsin\theta+W=500+\frac{2500}{3}=\frac{4000}3N"



effective area for pressure


"A=\pi r^2=3.14\times25\times10^{-4}=0.2m^2"



pressure at point A , B , and C is as following


"P_A=\frac{N_A}{A}=\frac{2000}{0.6}=3.33KPa"


"P_B=\frac{N_B}{A}=\frac{4000}{0.6}=6.66KPa"


"P_C=\frac{N_c}{A}=\frac{2000}{0.6}=3.33KPa"